OpenStudy (anonymous):

Choose the system of linear inequalities that represent the given scenario. Let a represent the number of adult tickets and let c represent the number of children tickets. There were up to 1,200 tickets available for the spring carnival. Adult tickets cost $6.00 and children’s tickets cost $3.50. How many tickets of each could be sold to make at least $6,450?

OpenStudy (anonymous):

Please help I missed this question.

OpenStudy (aum):

"up to 1,200 tickets" translates to "less than or equal to 1,200". If you add the number of adult and children tickets sold it should be \(\le 1200\). \(a + c \le 1200\) ------ (1) Each adult ticket costs $6. 'a' adult tickets will cost $6 * a. Each child's ticket costs $3.50. 'c' children tickets will cost $3.50 * c Total sales = 6a + 3.5c "make at least $6,450" means total sales must be "greater than equal to $6,450" \(6a + 3.5c \ge 6450\) ------ (2) Those are the two inequalities.

OpenStudy (anonymous):

Thanks :)

OpenStudy (aum):

You are welcome.